for preparation of 4% sucrose solution how much amount of sugar is added to 2 litres of water
Answer by JoelSchwartz(130)

x=total amount of sugar being added in liters
y=the number of liters of solution
y=x+2
.04y=x
.04(x+2)=x
.04x+.08=x
.08=x-.04x
.08=.96x
x=8/96